You roll a 6-sided die. What is the probability that you will roll either a 3 or a 2? P (3 or 2) = You flip a 2-sided coin. What is the probability that you will get either heads or tails? P (heads or tails) =
Answer #1

There are total 6 possible outcomes.

Hence,

P(3 or 2) = 2/6 = 1/3

There are total of two possible outcomes i.e. Heads or tails. Hence,

P(Heads or tails) = 2/2 = 1